OpenStudy (anonymous):

Given the equations \[9x+\frac{ 3 }{ 4 } y=6 ~and~ 2x+\frac{ 1 }{ 2 }y=9\], by what factor would you multiply the second equation to eliminate y and solve the system through linear combinations? A) \[-\frac{ 4 }{ 3 }\] B) \[−\frac{ 3 }{ 4 }\] C) \[−\frac{ 3 }{ 2 }\] D) \[−\frac{ 7 }{ 2 }\] @kelliegirl33 @Salmon

OpenStudy (anonymous):

9x + 3/4y = 6 2x + 1/2y = 9 hmmmm.....1/2 * -3/2 = - 3/4 you would multiply by -3/2, which would make the y in the second equation be -3/2.....and this will cause the y's to cancel out.

OpenStudy (anonymous):

oops..would make the y in the second equation be -3/4

OpenStudy (anonymous):

the answer is -3/2

OpenStudy (anonymous):

So the correct answer would be B?

OpenStudy (anonymous):

Oh, so C?

OpenStudy (anonymous):

yes...C is the answer

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

YW :)